A liquid flows through two capillary tubes A and B connected in series. The length and radius of B are twice those
of A. What is the ratio of the pressure difference across A to that across B?

Grade:

1 Answers

Abhishek Raj
6 years ago
According to continuity equation, dischargr will remain same in both tubes. i.e. Q1 = Q2According to poiseuilli` equationQ = ( P * pi * r^4)/ (8 * ω * L) Where r= radius of tube ω = viscosity of fluid L = length of tubeSolving Q1 = Q2 we getP1/P2 = 8 Ans.